OpenStudy (anonymous):

Solve each triangle using either the Law of Sines or the Law of Cosines. If no triangle exists, write “no solution.” Round your answers to the nearest tenth: a = 12, b = 14, c = 21 A = 51°, b = 41, c = 40 A = 23°, B = 55°, b = 9 A = 18°, a = 25, b = 18

OpenStudy (campbell_st):

Question 1 Uses the law of cosines \[cos(A) = \frac{b^2 +c^2-a^2}{2bc}\] once you have A you can use the law of sines to find a 2nd angle then the angle sum of a triangle for the 3rd angle. Question 2 Law of cosines \[a^2 = b^2 + c^2 - 2bc \cos(A)\] Question 3 law of sines \[\frac{a}{\sin(A)} = \frac{b}{\sin(B)} = \frac{c}{\sin(C)}\] Question 4 Law of Sine hope it helps... just substitute and evaluate

OpenStudy (anonymous):

i just need 3 and 4

OpenStudy (anonymous):

Cosine Law: c² = a² + b² – 2ab cos C Sine Law: Sine a / a = Sine b / b = sine c / c Question 1, you have a, b, and c. You can use the cosine law to solve for angle A, B, and C. Question 2: You can rewrite the cosine law to A² = B² + C² - 2BC cos A and solve for angle A. After you have the angle, you can use sine law to solve for angle B and C. Question 3: You can use Sine law to solve for A, and you can calculate angle C (sum of triangle) and thus length of C using the Cosine law Question 4: Similar to Question 3. Use Sine law to solve for angle B, and thus angle C, and then C using cosine